# Orthogonal projections of hypercubes

### v3

@@ VII.5 The hypercubic energy landscape of the three-ring system
Unlike the one-ring and two-ring systems, the structure of pathways is more complex, but we can still infer the state transition from the projection. For example, the pathway from the ground state $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\circ&\circ&\circ&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-3&0.9\end{bmatrix}^{\top}$) to $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-1.5&3.2\end{bmatrix}^{\top}$) is a transition with the correlated spin flips of spins located in the lower right hexagon (spins 3, 4, 7, and 11). Pathways parallel to this pathway involve the same spin flips. From the biplot basis on the top right of Fig. 10(a3), one can generally know that by following the angle of each biplot basis, the transitions involve spins on the top hexagon (spins 12, 13, 14, and 15) being flipped: the pathways are lines from the left top to the right bottom, which are parallel to the biplot basis of spins 12, 13, 14, and 15. Likewise, horizontal pathways are used when spins in the lower left hexagon (spins 1, 2, 5, and 8) are flipped, and vertical pathways are followed when spins in the inverted triangle at the center (spins 6, 9, and 10) are flipped. Similar to the one-ring and two-ring systems, the pathway is a collective flip of spins reflecting the interaction network. Notice that this projection of the hypercubic energy landscape [Fig. 10(a3)] is a partial Hamming projection: the PC1 score is equivalent to the Hamming distance from the ground state on the left side (around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$) or the right side (around $\begin{bmatrix}3&0\end{bmatrix}^{\top}$) but without considering spins 6, 9, and 10.
@@
To understand the relation between PC loading and the interaction network, in particular, the relation between PC loading and the frustration of the system, we color the loading of PC1 and PC2 in Figs. 10(b3) and 10(c3), following the angle of the biplot basis in Fig. 10(a3). We confirmed that the element-wise sign of PC1 [Fig. 10(b3)] corresponds to the ground state group containing six ground states, but the states of spins 6, 9, and 10 are not determined. These spins involve a frustrated cycle of the interaction network. PC2 loading in Fig. 10(c3) shows that spins 1, 2, 5, and 8 do not contribute, and those spins belong to the lower left ring. Moreover, PC2 loading of spins 6, 9, and 10 has slightly higher or lower magnitudes than the others, indicating the uncommon contribution of spins 6, 9, and 10, suggesting the frustration-related correlation.
@@
These insights are consistent with the interaction network with colored nodes in Fig. 10(d3): the PC loading captures the correlation arising from the interaction network. Nevertheless, spins 6 and 9 have exactly the same color: their biplot basis are the same, but it can be another combination of spins such as 6 and 10. This breaks the symmetry of the system—PC2 ignores the lower left hexagon of the interaction network even though the system has the same other structure—and we examine how the PC loading reflects the symmetry of the system.
VII.6 Fraction of explained variance
@@
VII.7 The hypercubic energy landscape of the three-ring system through PC3
The fraction of explained variance of the three-ring system [Fig. 11(a3)] leads us to visualize the hypercubic energy landscapes of three-ring101010 See the Supplemental Material [109] for hypercubic energy landscapes of two-ring systems using PC3. systems using PC3. In Fig. 12(a), the hypercubic energy landscape of three-ring systems, the same as Fig. 10(a3), is shown but by PC1 and PC3. We find that with PC1 and PC3, unlike the projection by PC1 and PC2, the shape of the hypercubic energy landscape looks similar to a parallelogram, not a hexagon. The biplot basis on the top right of Fig. 12(a) are only six arrows despite the system having 15 spins, indicating numerous overlaps of the biplot basis. To check which spins are overlapped, we show the loading of PC1 and PC3 in Figs. 12(b) and 12(c), and the interaction network in Fig. 12(d), coloring them by the angle of the biplot basis. PC3 loading [Fig. 12(c)], in particular, has a nonuniform magnitude of the elements; the magnitude of the elements of spins 1, 2, 5, 6, 8, and 9 are larger than the others. Those spins belong to the lower left hexagon of the interaction network in Fig. 12(d). The interaction network colored by the angle of the biplot basis in Fig. 12(d) illustrates how the biplot basis capture the correlation of spins. As indicated by PC3 loading [Fig. 12(c)], the spins on the lower left hexagon (spins 1, 2, 5, 6, 8, and 9) form the (anti)correlated group. The rest of the two rings form a large (anti)correlated group. PC3 loading creates the hypercubic energy landscape, emphasizing the lower left hexagon of the interaction network.
@@ VII.7 The hypercubic energy landscape of the three-ring system through PC3
Figure 12: Same as Figs. 10(a3), 10(b3), 10(c3), and 10(d3) but by different PCs. (a) The hypercubic energy landscape of the three-ring system by PC1 and PC3. (b) PC1 loading of (a) and the color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC3. (d) The Ising spin interaction networks of the three-ring systems, where the nodes are colored by the angle of the corresponding biplot basis in (a). (e–h) Same as (a–d) but by PC2 and PC3. Notice that in (a) and (e), several biplot basis are overlapped.
@@
We further investigate the hypercubic energy landscape of the three-ring system by PC2 and PC3 to understand the symmetry of the system. The hypercubic energy landscape of the three-ring system by PC2 and PC3 [Fig. 12(e)], looks like a hexagon, similar to the projection by PC1 and PC2. Nevertheless, the landscape is more symmetric and regular than that by PC1 and PC2. Around the origin, the connected six ground states also form a regular hexagon, but they are overlapping with the other group of six ground states, cf. Fig. 10(a3). The angles of the biplot basis are almost uniform (with some overlapping), and the lengths of them are nearly the same, resulting in a symmetric and regular hypercubic energy landscape; this projection is the same kind as the isometric projection, but that of a nine-dimensional hypercube.
@@
The loading of PC2 and PC3 is shown in Figs. 12(f) and 12(g), with the node color indicating the angle of the biplot basis. As we mentioned in Sec. VII.5, the PC2 loading emphasizes contributions from spins 6, 9, and 10 but ignores spins 1, 2, 5, and 8. On the contrary, PC3 loading has a higher magnitude for spins 1, 2, 5, and 8 in addition to spins 6 and 9, but ignores spin 10. Spins belonging to other hexagons contribute equally to the PC2 and PC3 loading. Therefore, as shown in the biplot basis of Fig. 12(a), the contributions by PC2 and PC3 are complementary, and the resulting hypercubic energy landscape shows the symmetry of the state space reflecting the interaction network.

### v4

@@ VII.6 The hypercubic energy landscape of the three-ring system
Unlike the one-ring and two-ring systems, the structure of pathways is more complex, but we can still infer the state transition from the projection. For example, the pathway from the ground state $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\circ&\circ&\circ&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-3&0.9\end{bmatrix}^{\top}$) to $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-1.5&3.2\end{bmatrix}^{\top}$) is a transition with the correlated spin flips of spins located in the lower right hexagon (spins 3, 4, 7, and 11). Pathways parallel to this pathway involve the same spin flips. From the biplot vectors on the top right of Fig. 10(a3), one can generally know that by following the angle of each biplot vector, the transitions involve spins on the top hexagon (spins 12, 13, 14, and 15) being flipped: the pathways are lines from the left top to the right bottom, which are parallel to the biplot vectors of spins 12, 13, 14, and 15. Likewise, horizontal pathways are used when spins in the lower left hexagon (spins 1, 2, 5, and 8) are flipped, and vertical pathways are followed when spins in the inverted triangle at the center (spins 6, 9, and 10) are flipped. Similar to the one-ring and two-ring systems, the pathway is a collective flip of spins reflecting the interaction network. Notice that this projection of the hypercubic energy landscape [Fig. 10(a3)] is a partial Hamming projection: the PC1 score is equivalent to the Hamming distance from the ground state on the left side (around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$) or the right side (around $\begin{bmatrix}3&0\end{bmatrix}^{\top}$) but without considering spins 6, 9, and 10.
@@
To understand the relation between PC loading and the interaction network, in particular, the relation between PC loading and the frustration of the system, we color the loading of PC1 and PC2 in Figs. 10(b3) and 10(c3), following the angle of the biplot vectors in Fig. 10(a3). We confirmed that the element-wise sign of PC1 [Fig. 10(b3)] corresponds to the ground state group containing six ground states, but the states of spins 6, 9, and 10 are not determined. These spins involve a frustrated cycle of the interaction network. PC2 loading in Fig. 10(c3) shows that spins 1, 2, 5, and 8 do not contribute, and those spins belong to the lower left ring. Moreover, PC2 loading of spins 6, 9, and 10 has slightly higher or lower magnitudes than the others, indicating the uncommon contribution of spins 6, 9, and 10, suggesting the frustration-related correlation.
@@
These insights are consistent with the interaction network with colored nodes in Fig. 10(d3): the PC loading captures the correlation arising from the interaction network. Nevertheless, spins 6 and 9 have exactly the same color: their biplot vectors are the same, but it can be another combination of spins such as 6 and 10. This breaks the symmetry of the system—PC2 ignores the lower left hexagon of the interaction network even though the system has the same other structure—and we examine how the PC loading reflects the symmetry of the system.
VII.7 Fraction of explained variance
@@
Figure 12: Same as Figs. 10(a3), 10(b3), 10(c3), and 10(d3) but by different PCs. (a) The hypercubic energy landscape of the three-ring system by PC1 and PC3. (b) PC1 loading of (a) and the color of the bar matches the color of the corresponding biplot vectors in (a). (c) Same as (b) but for PC3. (d) The Ising spin interaction networks of the three-ring systems, where the nodes are colored by the angle of the corresponding biplot vectors in (a). (e)–(h) Same as (a)–(d) but by PC2 and PC3. Notice that in (a) and (e), several biplot vectors are overlapped.
VII.8 The hypercubic energy landscape of the three-ring system through PC3
The fraction of explained variance of the three-ring system [Fig. 11(a3)] leads us to visualize the hypercubic energy landscapes of three-ring111111 See the Supplemental Material [16] for hypercubic energy landscapes of two-ring systems using PC3. systems using PC3. In Fig. 12(a), the hypercubic energy landscape of three-ring systems, the same as Fig. 10(a3), is shown but by PC1 and PC3. We find that with PC1 and PC3, unlike the projection by PC1 and PC2, the shape of the hypercubic energy landscape looks similar to a parallelogram, not a hexagon. The biplot vectors on the top right of Fig. 12(a) are only six arrows despite the system having 15 spins, indicating numerous overlaps of the biplot vectors. To check which spins are overlapped, we show the loading of PC1 and PC3 in Figs. 12(b) and 12(c), and the interaction network in Fig. 12(d), coloring them by the angles of the biplot vectors. PC3 loading [Fig. 12(c)], in particular, has a nonuniform magnitude of the elements; the magnitude of the elements of spins 1, 2, 5, 6, 8, and 9 are larger than the others. Those spins belong to the lower left hexagon of the interaction network in Fig. 12(d). The interaction network colored by the angles of the biplot vectors in Fig. 12(d) illustrates how the biplot vectors capture the correlation of spins. As indicated by PC3 loading [Fig. 12(c)], the spins on the lower left hexagon (spins 1, 2, 5, 6, 8, and 9) form the (anti)correlated group. The rest of the two rings form a large (anti)correlated group. PC3 loading creates the hypercubic energy landscape, emphasizing the lower left hexagon of the interaction network.
@@ VII.8 The hypercubic energy landscape of the three-ring system through PC3
We further investigate the hypercubic energy landscape of the three-ring system by PC2 and PC3 to understand the symmetry of the system. The hypercubic energy landscape of the three-ring system by PC2 and PC3 [Fig. 12(e)], looks like a hexagon, similar to the projection by PC1 and PC2. Nevertheless, the landscape is more symmetric and regular than that by PC1 and PC2. Around the origin, the connected six ground states also form a regular hexagon, but they are overlapping with the other group of six ground states, cf. Fig. 10(a3). The angles of the biplot vectors are almost uniform (with some overlapping), and the lengths of them are nearly the same, resulting in a symmetric and regular hypercubic energy landscape; this projection is the same kind as the isometric projection, but that of a nine-dimensional hypercube.
@@
The loading of PC2 and PC3 is shown in Figs. 12(f) and 12(g), with the node color indicating the angle of the biplot vectors. As we mentioned in Sec. VII.6, the PC2 loading emphasizes contributions from spins 6, 9, and 10 but ignores spins 1, 2, 5, and 8. On the contrary, PC3 loading has a higher magnitude for spins 1, 2, 5, and 8 in addition to spins 6 and 9, but ignores spin 10. Spins belonging to other hexagons contribute equally to the PC2 and PC3 loading. Therefore, as shown in the biplot vectors of Fig. 12(a), the contributions by PC2 and PC3 are complementary, and the resulting hypercubic energy landscape shows the symmetry of the state space reflecting the interaction network.